What is the sum of 2+4+6+8+10+12 using the formula for the partial sum of arithmetic series?
Answer by jim_thompson5910(35256)

S = (n*(a1 + an))/2

S = (6*(2+12))/2

S = 6*14/2

S = 84/2

S = 42

So 2+4+6+8+10+12 = 42